# Phase Matching Free Sensing with Undetected Light Using a Nonlinear Metasurface

### v1

@@ Abstract
In this letter, we report classical sensing with undetected light using octave spanning stimulated four-wave mixing from a plasmonic metasurface. The bidirectional nonlinear scattering due to inherent reflections from such thin nonlinear materials modifies their operation within a nonlinear interferometer. The theoretical model for visibility accounting for such bidirectionality as well as pulsed illumination accurately predicts visibility in the system as a function of transmission in the near-infrared seed (idler) arm. Spectrally resolving the visible signal emission evaluates the total dispersion within the interferometer, highlighting the prospect of ultrafast sensing with undetected photons.
@@ Introduction
Nonlinear interferometry allows phase and amplitude information to be passed between light that only interacts with an object and light that is only detected.1, 2, 3, 4 The interferometric technique uses correlated photon beams created in a nonlinear material, either via difference frequency generation3 or parametric fluorescence1. When implemented with non-degenerate beams, there are wide-reaching applications where passing information between wavelengths is desirable; for example, where the technical and fundamental limitations of mid-infrared (MIR) detectors can be circumvented with silicon-based sensors5, 6, 7 for label-free biological imaging,8, 9 environmental monitoring,10 and medical imaging.11, 12
@@
Recent research has predominantly focused on the quantum mechanical variant of nonlinear interferometry; where spontaneous parametric down conversion (SPDC), typically created in a bulk nonlinear crystal (NLC), interferes by induced coherence without induced emission.13, 1, 9, 14, 15, 16 Here, indistinguishable photon pairs are generated in two paths of an interferometer such that they occupy two overlapping beam paths, labelled here as the high energy signal and the low energy idler. Induced coherence without induced emission occurs due to the lack of "which path" information; an object placed in one of the idler paths, for example, introduces path knowledge that destroys interference. What differentiates quantum nonlinear interferometry from other quantum sensing techniques, such as ghost imaging17, is that the sensing information can be found by only detecting photons that never interacted with the object. Classical nonlinear interferometry achieves the same effect, but via stimulating the parametric fluorescence; for example, stimulating the idler while detecting the signal.4, 3, 18 The advantages of the classical variant over the quantum mechanical are improved signal-to-noise ratio and greater coherence of the bi-photon state3, 19.
@@
Quantum nonlinear interferometry research has resulted in low cost and portable16 near-infrared (NIR) imaging systems at video rate14, 20 and MIR operation.9, 15 While the operating frequency range is set by the properties of the more common NLCs, currently achieved at up to $\sim$4 µm with 800 nm detection,15 more exotic NLCs are being explored for operation up to 10 µm.21, 22, 23 However, crystals of order millimetres in length are required to generate sufficient levels of nonlinear emission through phase matching techniques. Alternative media that promise range extension are nonlinear metasurfaces—subwavelength thick, nanostructured arrays that enhance nonlinear processes by supporting electromagnetic resonances with high field localisation.24, 25, 26, 27 Metasurfaces can be tailor made for operation at targeted frequencies and their subwavelength thickness relaxes any phase matching and absorption constraints of the resonant nonlinearity.28, 29 While metasurfaces outperform unpatterned films of the same thickness and material in regards to spontaneous photon pair emission rates,24, 26 they are currently insufficient for implementing quantum nonlinear interferometry. However, by stimulating emission—and therefore operating in the classical regime—metasurfaces become suitable nonlinear sources.
@@
Here, we demonstrate classical nonlinear interferometry to achieve phase matching free sensing with undetected photons using a resonant gold plasmonic nonlinear metasurface, where four-wave mixing (FWM) is stimulated using femtosecond pulsed pump and seed illumination. In our case, the seed beam stimulates the interferometer’s NIR idler mode to produce FWM in the visible signal mode. Amplitude and phase information of one path of the idler is observed in the visibility and phase of the signal mode interference. Operated as a folded Michelson interferometer, we study the influence of pump, seed and idler reflections combined with the bidirectional FWM emission of the metasurface. Unlike conventional nonlinear crystals, metasurface reflections cannot be eliminated due to their resonant nature. Despite these reflections and the high order nonlinearity involved, we find that the homodyne advantage leads to a signal interference visibility of over 50%. We also find signal interference independent of pump-seed delay and discuss how this can be used for ultrafast sensing with undetected light. Although experimentally implemented with a metasurface, the theoretical framework is appropriate for any thin-film nonlinear source.
Results and discussion
0.0.1 Nonlinear interferometry with a subwavelength source
@@
In the Michelson-style interferometer, Fig. 1(a), FWM (signal) is generated on two pump and seed passes through the nonlinear metasurface, shown in Fig. 1(b). Pulses at 835 nm (pump, red) and 1500 nm (seed, magenta) are overlapped in transverse space at a 1 µm short-pass dichroic mirror, and in time using a piezo-controlled delay line in the seed path. Both beams are expanded to $\sim$10 mm in diameter and focused using a $20$ cm focal length lens to illuminate a $75\times 75$ µm2 metasurface across an areal overlap of 21 µm diameter. The metasurface is composed of gold antennas that are doubly resonant at the pump and seed wavelengths leading to enhanced electric field overlap at the incident wavelengths, generating FWM emission at 580 nm (signal, yellow) and 1500 nm (idler, magenta). Both the linear and nonlinear responses of the structures have been previously reported.30, 31 The subwavelength thickness of the nonlinear metasurface relaxes out-of-plane phase matching, resulting in FWM (signal) in both directions on each pass, illustrated in Fig. 1(c, d). Combined with reflections of the various wavelengths from the metasurface, the nonlinear interference is distinct from conventional nonlinear interferometry.
@@
Figure 1: The interferometer system, metasurface and interference (a) The Michelson interferometer setup operating in the far field. The 835 nm pump (red) and 1500 nm seed (idler, magenta) pulsed lasers stimulate 580 nm FWM (signal, yellow). (b) SEM image the nonlinear metasurface comprised of two gold discs and a gold bar antenna per unit cell. The scale bar in 500 nm. (c, d) Graphical illustration of the bidirectional FWM (signal) emission at the first and second pass, respectively. The optical axes have been angularly displaced for clarity and do not indicate amplitude. (e) Temporal positions of FWM (signal) pulses created at each pass. The dashed lines represent $E_{f1+}$ and $E_{f2-}$. Distance $d$ is from the metasurface to curved mirror 1. (f, g, h) Intensity of FWM (signal) at the camera with no beam blocks in the interferometer arms, when beams are blocked at curved mirror 1, and when the seed (idler) is blocked at curved mirror 2, respectively. (i) Intensity at the camera along the dashed lines on (f, g, h).
@@
We define the frequency domain pump and seed fields on the first pass of the metasurface as ${{\bf E}_{p1}=E_{p}(\omega)\hat{x}}$ and ${{\bf E}_{s1}=E_{s}(\omega)\hat{y}}$, respectively, where polarisation is relative to the axes marked on Fig. 1(b) . FWM (signal) is generated as $\omega_{f}=2\omega_{p}-\omega_{s}$. After the metasurface, the seed (idler) is separated from the pump and FWM (signal) at a 1 µm long-pass dichroic mirror, and a pair of identical curved mirrors with $20$ cm radius of curvature reflect all wavelengths back along their respective paths to the metasurface. On the second pass, the returning pump field at the metasurface is ${\bf E}_{p2}=\sqrt{T_{p}}{\bf E}_{p1}e^{2ik_{p}d}$, where $T_{p}$ is the metasurface pump transmission and $d$ is the optical path length of the pump and FWM (signal) arm of the interferometer. Meanwhile the returning seed field is ${\bf E}_{s2}=\sqrt{T_{s}T}e^{i\gamma}{\bf E}_{s1}e^{2ik_{s}d^{\prime}}$, where $T_{s}$ is the seed transmission of the metasurface, $T$ is the transmission coefficient and $\gamma$ the phase shift of the object to be sensed, and $d^{\prime}$ is the optical path length of the interferometer’s seed (idler) arm. FWM (signal) emission at each pass occurs in both $+$ and -- directions. The FWM field amplitude on the first pass is $\sigma_{\pm}(\omega)=\alpha_{\pm}E_{p}^{2}E_{s}^{*}$, where $\alpha_{\pm}$ indicates the strength of the nonlinear process, which is related to the nonlinear $\chi^{(3)}_{\mu xxy}$ tensor elements for the process. FWM (signal) on the second pass is reduced by the transmission and phase at each wavelength. We thus identify four FWM (signal) amplitudes at the metasurface per incident pump and seed pulse, illustrated and labelled in Figure 1 (c, d), and evaluate their resulting amplitude and phase at the camera to be
$E_{f1-}=\sigma_{-}$ (1)
@@
where $T_{f}$ is the metasurface FWM (signal) transmission, $2k_{f}(d^{\prime}-d)=\psi$, is the phase difference between the interferometer’s arms. The normally incident and expanded pump and seed beams ensure that the nonlinear emission is nearly in-phase across the metasurface, equivalent to matching of the in-plane momenta of the various fields, so that the longitudinal momentum is $2k_{p}-k_{s}=k_{f}$. The FWM (signal) is isolated from the pump and seed and directed towards a silicon camera using a $650$ nm long-pass dichroic mirror. The camera thus integrates over three FWM pulses per input pump and seed pulse with relative time delays of 00, $2d/c$, $4d/c$, as illustrated in Fig. 1(e). Since the delay between distinct FWM pulses, $2d/c$, is greater than the FWM (signal) pulse duration, no interference occurs between them and the camera measures a time-averaged intensity.
@@
The third term in Eq. (5) is an interference of the forward scattered FWM (signal) from the second pass, $E_{f2+}$, with the forward scattered FWM (signal) from the first pass reflected by the metasurface, $\sqrt{R_{f}}E_{f1+}$ where $R_{f}$ is the metasurface FWM reflectivity. Curved mirror 1 is connected to a second piezo-controlled stage, allowing the phase difference, $\psi$, to be varied. The detected signal normalised to the forward FWM emission signal, $\sigma_{+}^{2}$, is
@@
where $\kappa=T_{p}^{2}T_{s}$ is the ratio of total FWM (signal) from the two passes, and $\eta=\sigma_{-}/\sigma_{+}$ is the ratio of bidirectional emission. A single frame of the FWM interference at the camera can be seen in Fig. 1(f). When a beam block is placed in front of curved mirror 1, the forward emitted first pass, $E_{f1+}$, is no longer detected and the pump beam does not generate FWM (signal) at the second pass, $E_{f2\pm}=0$, so no interference is observed in Fig. 1(g), and the camera measures $\sigma_{-}^{2}$. When the beam block is placed in front of curved mirror 2, the seed no longer stimulates second pass emission, $E_{f2\pm}=0$, and interference is destroyed in Fig. 1(h), so the camera records $\sigma_{-}^{2}+(1+R_{f})T_{f}\sigma_{+}^{2}$. Line plots in Fig. 1(i) clarify interference and its subsequent breakdown in Fig. 1(f-h). A video of the interference fringes as $\psi$ is varied can be found in the Supporting Information (SI). By characterising the linear transmissivity and reflectivity of each field, shown in Table I of the SI, we can evaluate the theoretical performance of the metasurface nonlinear interferometer. The expected visibility is $V=(N_{max}-N_{min})/(N_{max}+N_{min})$ such that
@@
At $T=100\%$, the theoretical visibility predicts a maximum of 55.3 $\pm$ 0.5%, when applying $\eta$ extracted from the images in Fig. 1(g) and (h). For reference, a conventional nonlinear crystal with ideal anti-reflection coatings placed within this interferometer would only have two interfering FWM (signal) fields, $D_{0}=\sigma_{0}^{2}|E_{f1+}+E_{f2-}|^{2}$, which for collinear phase-matched emission, implies $\sigma_{\pm}=\sigma_{0}$, such that, $N_{0}=1+T+2\sqrt{T}\cos(\gamma+\psi)$, with a maximum visibility, $V_{0}=2\sqrt{T}/(1+T)\mapsto 1$ in the case of no object. A more detailed derivation for the visibility can be found in Section I of the SI.

### v2

# Phase Matching Free Sensing with Undetected Light Using a Nonlinear Thin-Film Metasurface

@@ Abstract
In this article, we report classical sensing with undetected light using octave spanning stimulated four-wave mixing from a plasmonic metasurface. The bidirectional nonlinear scattering due to inherent reflections from such thin nonlinear materials modifies their operation within a nonlinear interferometer. The theoretical model for visibility accounting for such bidirectionality as well as pulsed illumination accurately predicts visibility in the system as a function of transmission in the near-infrared seed (idler) arm. Spectrally resolving the visible signal emission evaluates the total dispersion within the interferometer, highlighting the prospect of ultrafast sensing with undetected photons.
@@ Introduction
Nonlinear interferometry allows phase and amplitude information to be passed between light that only interacts with an object and light that is only detected.Lemos2014_QUIP_OG, Lahiri2015_QUIP_Theory, Cardoso_2018_ClassicalIUL, Shapiro2015_ClassicalIUL_Theory The interferometric technique uses correlated photon beams created in a nonlinear material, either via difference frequency generationCardoso_2018_ClassicalIUL or parametric fluorescenceLemos2014_QUIP_OG. When implemented with non-degenerate beams, there are wide-reaching applications where passing information between wavelengths is desirable; for example, where the technical and fundamental limitations of mid-infrared (MIR) detectors can be circumvented with silicon-based sensorsRogalski_2016_IRdetector_challenges, Lindner2020_FTIR_UP_low-gain, Lindner2023_MIRsensing_UP_SiDetect for label-free biological imaging,Greaves2023_SNOM_intracell_labelFree, Kviatkovsky2020_QIUP_MIR environmental monitoring,Myers2009_MIRsensing_OrganPollutant and medical imaging.Amrania2018_Digistain, Contreras-Rozo2023_MIRspecDiabetes
@@
Recent research has predominantly focused on the quantum mechanical variant of nonlinear interferometry; where spontaneous parametric down conversion (SPDC), typically created in a bulk nonlinear crystal (NLC), interferes by induced coherence without induced emission.Wang1991_InducedCoherenceWithout, Lemos2014_QUIP_OG, Kviatkovsky2020_QIUP_MIR, Basset2021_videoRateQIUP, Kviatkovsky2022_QIUP_MIR_Microscopy, Pearce2023_compact Here, indistinguishable photon pairs are generated in two paths of an interferometer such that they occupy two overlapping beam paths, labelled here as the high energy signal and the low energy idler. Induced coherence without induced emission occurs due to the lack of "which path" information; an object placed in one of the idler paths, for example, introduces path knowledge that destroys interference. What differentiates quantum nonlinear interferometry from other quantum sensing techniques, such as ghost imagingErkmanShapiro_2010_GhostReview, is that the sensing information can be found by only detecting photons that never interacted with the object.
@@
Quantum nonlinear interferometry research has resulted in low cost and portablePearce2023_compact near-infrared (NIR) imaging systems at video rateBasset2021_videoRateQIUP, Pearce2024_VideoRate_QIUP and MIR operation.Kviatkovsky2020_QIUP_MIR, Kviatkovsky2022_QIUP_MIR_Microscopy While the operating frequency range is set by the properties of the more common NLCs, currently achieved at up to 4 µm with 800 nm detection,Kviatkovsky2022_QIUP_MIR_Microscopy more exotic NLCs are being explored for operation up to 10 µm.Kumar2021_AgGaS2_photonPair, Paterova2022_Qspec_AgGaS2_8um, Mukaii2022_QFTIR_AgGaS2 However, crystals of order millimetres in length are required to generate sufficient levels of nonlinear emission through phase matching techniques. Alternative media that promise range extension are nonlinear metasurfaces; subwavelength thick, nanostructured arrays that enhance nonlinear processes by supporting electromagnetic resonances with high field localisation.Satiago-Cruz2021_LN_metasurface, Santiago-Cruz2022_GaAs_BIC_metasurface, Son2023_SPDC_BICbidirectional_MS, Zhang2024_photonIndistinguish_SiMetasurface Metasurfaces can be tailor-made for operation at targeted frequencies and their subwavelength thickness relaxes any phase matching and absorption constraints of the resonant nonlinearity.Schulz2024_MetasurfaceRoadMap, Oulton2022_NPnews&views_thinNL While metasurfaces outperform unpatterned films of the same thickness and material in regards to spontaneous photon pair emission rates,Satiago-Cruz2021_LN_metasurface, Son2023_SPDC_BICbidirectional_MS they are currently insufficient for implementing quantum nonlinear interferometry. However, by stimulating emission – and therefore operating in the classical regime – metasurfaces become suitable nonlinear sources.
Classical nonlinear interferometry achieves the same effect as the quantum implementation, passing amplitude and phase information across wavelengths, but via stimulating the parametric fluorescence; for example, stimulating the idler while detecting the signal.Shapiro2015_ClassicalIUL_Theory, Cardoso_2018_ClassicalIUL, Huang2023_fibreStimFWMimaging The advantages of the classical variant over the quantum mechanical are an improved signal-to-noise ratio and greater coherence of the bi-photon state.Cardoso_2018_ClassicalIUL, Kolobov_2017_ControllingInducedConherence Upconversion sensing also relies on stimulated nonlinear emission;Johnson2012_MIRfs_upcon, Ashik2019_MIRfs_upcon however, classical nonlinear interferometry distinguishes itself with phase sensitivity, stronger robustness to noise, enhanced sensitivity at low amplitude, and immunity to incoherent backgrounds.
@@
Here, we demonstrate classical nonlinear interferometry to achieve phase matching free sensing with undetected photons using a resonant gold plasmonic nonlinear metasurface, where four-wave mixing (FWM) is stimulated using femtosecond pulsed pump and seed illumination. In our case, the seed beam stimulates the interferometer’s NIR idler mode to produce FWM in the visible signal mode. Amplitude and phase information of one path of the idler is observed in the visibility and phase of the signal mode interference. Operated as a folded Michelson interferometer, we study the influence of pump, seed and idler reflections combined with the bidirectional FWM emission of the metasurface. Unlike conventional nonlinear crystals, metasurface reflections cannot be eliminated due to their resonant nature. Despite these reflections and the high-order nonlinearity involved, we find that the homodyne advantage leads to a signal interference visibility of over 50%. We also find signal interference independent of pump-seed delay outside the interferometer and discuss how this can be used for ultrafast sensing with undetected light. Although experimentally implemented with a metasurface, the theoretical framework is appropriate for any thin-film nonlinear source.
Results & discussion
0.0.1 Nonlinear interferometry with a thin-film source
@@
In the Michelson-style interferometer, Fig. 1(a), FWM (signal) is generated on two pump and seed passes through the nonlinear metasurface, shown in Fig. 1(b). Pulses at 835 nm (pump, red) and 1500 nm (seed, magenta) are overlapped in transverse space at a 1 µm short-pass dichroic mirror, and in time using a piezo-controlled delay line in the seed path. Both beams are expanded to $\sim$10 mm in diameter and focused using a $20$ cm focal length lens to illuminate a $75\times 75$ µm2 metasurface across an areal overlap of 23.4 µm diameter. The peak powers delivered by the pump and seed fields are 14.5 kW and 80.0 W, respectively. The metasurface is composed of gold antennas that are doubly resonant at the pump and seed wavelengths leading to enhanced electric field overlap at the incident wavelengths, generating FWM emission at 580 nm (signal, yellow) and 1500 nm (idler, magenta). The subwavelength thickness of the nonlinear metasurface relaxes out-of-plane phase matching, resulting in FWM (signal) in both directions on each pass, illustrated in Fig. 1(c, d). Combined with reflections of the various wavelengths from the metasurface, the nonlinear interference is therefore distinct from conventional nonlinear interferometry. From a single pass through the metasurface, the peak power of FWM (signal) emission is $4.81$ nW across a $1.17$ mm diameter spot at the camera. The photon conversion efficiency relative to the seed beam is estimated to be $1.18\times 10^{-11}$. When the seed beam is blocked, no FWM (signal) is observed at the camera using typical settings of $800$ ms exposure time and $600\times$ EM gain. Both the linear and nonlinear responses of the structures have been studied in greater detail in previous literature.Gennaro2016_BarandDiscsSHG, Gennaro2018_barDiscs_PulseChar, Yang2025_StimTomography_B&D
@@
Figure 1: The interferometer system, metasurface and interference (a) The Michelson interferometer setup operating in the far field. The 835 nm pump (red) and 1500 nm seed (idler, magenta) pulsed lasers stimulate 580 nm FWM (signal, yellow). (b) SEM image the nonlinear metasurface comprised of two gold discs and a gold bar antenna per unit cell. The scale bar is 500 nm. (c, d) Graphical illustration of the bidirectional FWM (signal) emission at the first and second pass, respectively. The optical axes have been angularly displaced for clarity and do not indicate amplitude. (e) Temporal positions of FWM (signal) pulses created at each pass. The dashed lines represent $E_{f1+}$ and $E_{f2-}$. Distance $d$ is from the metasurface to curved mirror 1. (f, g, h) Intensity of FWM (signal) at the camera with no beam blocks in the interferometer arms, when beams are blocked at curved mirror 1, and when the seed (idler) is blocked at curved mirror 2, respectively. Scale bars are 2 mm. (i) Intensity at the camera along the dashed lines on (f, g, h).
@@
We define the frequency domain pump and seed fields on the first pass of the metasurface as $^{{\bf E}_{p1}=E_{p}(\omega)e^{i\phi_{p}(\omega)}\hat{x}}$ and $^{{\bf E}_{s1}=E_{s}(\omega)e^{i\phi_{s}(\omega)}\hat{y}}$, respectively, where polarisation is relative to the axes marked on Fig. 1(b). FWM (signal) is generated as $\omega_{f}=2\omega_{p}-\omega_{s}$. After the metasurface, the seed (idler) is separated from the pump and FWM (signal) at a 1 µm long-pass dichroic mirror, and a pair of identical curved mirrors with $20$ cm radius of curvature reflect all wavelengths back along their respective paths to the metasurface. On the second pass, the returning pump field at the metasurface is ${\bf E}_{p2}=\sqrt{T_{p}}{\bf E}_{p1}e^{2ik_{p}d}$, where $T_{p}$ is the metasurface pump transmission and $d$ is the optical path length of the pump and FWM (signal) arm of the interferometer. Meanwhile, the returning seed field is ${\bf E}_{s2}=\sqrt{T_{s}T}e^{i\gamma}{\bf E}_{s1}e^{2ik_{s}d^{\prime}}$, where $T_{s}$ is the seed transmission of the metasurface, $T$ is the transmission coefficient and $\gamma$ is the phase shift of the object to be sensed, and $d^{\prime}$ is the optical path length of the interferometer’s seed (idler) arm. FWM (signal) emission at each pass occurs in both ++ and - directions. The FWM field amplitude on the first pass is $\sigma_{\pm}(\omega)=\alpha_{\pm}E_{p}^{2}E_{s}^{*}$, where $\alpha_{\pm}$ indicates the strength of the nonlinear process, which is related to the nonlinear $\chi^{(3)}_{\mu xxy}$ tensor elements for the process. The spontaneous emission is assumed to be negligible in comparison to the stimulated emission. FWM (signal) on the second pass is reduced by the transmission and phase at each wavelength. We thus identify four FWM (signal) amplitudes at the metasurface per incident pump and seed pulse, illustrated and labelled in Fig. 1(c, d), and evaluate their resulting amplitude and phase at the camera to be
$E_{f1-}=\sigma_{-}=\alpha_{-}E_{p1}^{2}E_{s1}^{*}e^{ik_{f}l}$ (1)
@@
where $T_{f}$ is the metasurface FWM (signal) transmission, $2k_{f}(d^{\prime}-d)=\psi$, is the phase difference between the interferometer’s arms. The normally incident and expanded pump and seed beams ensure that the nonlinear emission is nearly in-phase across the metasurface, which is equivalent to matching the in-plane momenta of the various fields, such that the longitudinal momentum is $2k_{p}-k_{s}=k_{f}$. The FWM (signal) is isolated from the pump and seed and directed towards a silicon camera using a $650$ nm long-pass dichroic mirror. The camera thus integrates over three FWM pulses per input pump and seed pulse with relative time delays of 0, $2d/c$ and $4d/c$, as illustrated in Fig. 1(e). Since the delay between distinct FWM pulses, $2d/c$, is greater than the FWM (signal) pulse duration, no interference occurs between them, and the camera measures a time-averaged intensity
@@
The third term in Eq. (5) is an interference of the forward scattered FWM (signal) from the second pass, $E_{f2+}$, with the forward scattered FWM (signal) from the first pass reflected by the metasurface, $\sqrt{R_{f}}E_{f1+}$, where $R_{f}$ is the metasurface FWM reflectivity. Linear reflections of the pump and seed beams by the metasurface have been omitted due to nonlinear scaling of FWM emission. Curved mirror 1 is connected to a second piezo-controlled stage, allowing the phase difference, $\psi$, to be varied. The detected signal normalised to the forward emitted FWM (signal) emission signal, $\sigma_{+}^{2}$, is
@@
where $\kappa=T_{p}^{2}T_{s}$ is the ratio of total FWM (signal) from the two passes and $\eta=\sigma_{-}/\sigma_{+}$ is the ratio of bidirectional emission. A single frame of the FWM interference at the camera can be seen in Fig. 1(f). When a beam block is placed in front of curved mirror 1, the forward emitted first pass, $E_{f1+}$, is no longer detected and the pump beam does not generate FWM (signal) at the second pass, $E_{f2\pm}=0$, so no interference is observed in Fig. 1(g), and the camera measures $\sigma_{-}^{2}$. When the beam block is placed in front of curved mirror 2, the seed no longer stimulates second pass emission, $E_{f2\pm}=0$, and interference is destroyed in Fig. 1(h), so the camera records $\sigma_{-}^{2}+(1+R_{f})T_{f}\sigma_{+}^{2}$. Line plots in Fig. 1(i) clarify interference and its subsequent breakdown in Fig. 1(f-h). A video of the interference fringes as $\psi$ is varied can be found in the Supporting Information (SI). By characterising the linear transmissivity and reflectivity of each field, shown in Table I of the SI, we can evaluate the theoretical performance of the metasurface nonlinear interferometer. The expected visibility is $V=(N_{max}-N_{min})/(N_{max}+N_{min})$ such that
@@
At $T=100\%$, the theoretical visibility predicts a maximum of 55.3 $\pm$ 0.5% when applying $\eta$ extracted from the images in Fig. 1(g) and (h). For reference, a conventional nonlinear crystal with ideal anti-reflection coatings placed within this interferometer would only have two interfering FWM (signal) fields, $D_{0}=\sigma_{0}^{2}|E_{f1+}+E_{f2-}|^{2}$; which for collinear phase-matched emission, implies $\sigma_{\pm}=\sigma_{0}$, such that $N_{0}=1+T+2\sqrt{T}\cos(\gamma+\psi)$, with a maximum visibility, $V_{0}=2\sqrt{T}/(1+T)\mapsto 1$ in the case of no object. A more detailed derivation of the visibility can be found in Section SII of the SI. We note that the reduction in the maximum visibility when using a metasurface source is predominantly due to the $E_{f1-}$ field. The FWM (signal) fields can be directionally filtered in a Mach-Zehnder interferometer to achieve a maximum visibility that is theoretically equivalent to conventional nonlinear interferometers. Alternatively, tilting the metasurface to misalign nonlinear reflections also maximises visibility while simultaneously allowing the nonlinear emission to be monitored. These configurations are discussed theoretically in Sections SIII and SIV of the SI.